OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2).

OpenStudy (anonymous):

Help me with this please?

OpenStudy (anonymous):

y = mx +b slope:\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

can u find the slope now?

OpenStudy (anonymous):

After you find the slope "m", as @surdawi explained very well, then take either of those points of form (x1, y1) and use:\[y - y _{1} = m(x - x _{1})\] for substitution for x1 and y1. The "x" and the "y" stay as is as variables.

OpenStudy (anonymous):

So y=1/2x + 4?

OpenStudy (anonymous):

Yes! You got that right! Good job!